You want to determine whether the reasons given by workers for continuing their education is related to job type. In the study, you randomly collect the data shown in the contingency table. At ?=0.10, can you conclude that the reason and type of worker are dependent? Complete parts (a) through (d) Reason of worker Professional Personal Both echnical ther 35 40 43 37 22 31 (a) Identify the claim and state the null and alternative hypotheses H Reasons are H Reasons are The (b) Determine the degrees of freedom, find the critical value, and identify the rejection region What are the degrees of freedom? d.f. Find the critical value V the type of worker V the type of worker ?| is the claim. ?=? Round to three decimal places as needed.)
Answer #1

a)Ho:reasons are independent to the type of gender.

Ha: reasons are dependent to the type of gender.

b)

df =(row-1)*(column-1)=(2-1)*(3-1)=2

crtiical value =4.605

option C is correct

c)

applying chi square test:

OBS professional personal both Total
technical 35 43 37 115
other 40 22 31 93
Total 75 65 68 208
Ei=?row*?column/?total professional personal both Total
technical 41.466 35.938 37.596 115
other 33.534 29.063 30.404 93
Total 75 65 68 208
=(Oi-Ei)2/Ei professional personal both Total
technical 1.008 1.388 0.009 2.406
other 1.247 1.716 0.012 2.975
Total 2.255 3.104 0.021 5.381

test statistic =5.381

d) option B is correct